OpenStudy (anonymous):

Nine more than half the number n is no more than -8. Find n. Write an inequality. Then solve.

OpenStudy (anonymous):

1/2n + 9 <= -8 can you solve that ?

OpenStudy (anonymous):

1. subtract 9 from both sides 2. divide both sides by 1/2

OpenStudy (anonymous):

Whats negative one divided by one half?

OpenStudy (anonymous):

1/2n + 9 <= -8 --- subtract 9 1/2n <= -8 - 9 1/2n <= - 17 -- divide by 1/2 n <= -17/(1/2) n <= -17 * 2/1 n <= - 34

OpenStudy (anonymous):

do you understand what I did ? Do you have any questions ?

OpenStudy (acxbox22):

-1/0.5 = -2

OpenStudy (anonymous):

you do not even use -1 divide by 1/2 in this problem